OpenStudy (nekoaisu):

Solve by applying the zero product property. 8m(m+3)=2m-15 Find the solution set.

OpenStudy (tkhunny):

You'll need to get just a zero on one side.

OpenStudy (nekoaisu):

@tkhunny Could you please explain how, or give me an example?

OpenStudy (tkhunny):

Ever see one of these? \(ax^{2} + bx + c = 0\) It's a quadratic equation. It MUST have zero alone, by itself, on one side. You do this by addition and subtraction. 8m(m+3) = 2m - 15 Subtract 2m from each side 8m(m+3) - 2m = 2m - 15 - 2m 8m(m+3) - 2m = 2m - 2m - 15 8m(m+3) - 2m = 0 - 15 8m(m+3) - 2m = - 15 You may wish to expand the left hand side using the distributive property. 8m^2 + 24m - 2m = - 15 Sweet. Some "like" terms we can combine. 8m^2 + 22m = - 15 Okay, now you take care of the -15.

OpenStudy (nekoaisu):

Is this right? \[8m^2+22m+15\]

OpenStudy (nekoaisu):

@tkhunny

Directrix (directrix):

@NekoAisu Yes, if you write = 0 at the end. This is an equation. 8 m ^2 + 22 m +15 = 0

OpenStudy (nekoaisu):

@Directrix Thank you, but what do I need to do now? I need to find the solution set. Do I try to factor it?

Directrix (directrix):

Yes. Factor this: 8 m ^2 + 22 m +15 = 0

Directrix (directrix):

(4m + 5) * ( ? + ?) = 0 I did one factor. You do the other, okay?

OpenStudy (nekoaisu):

(2m+3)? @Directrix

Directrix (directrix):

(4m + 5) * ( 2m + 3) = 0 Zero Product Property 4m + 5 = 0 OR 2m + 3 = 0 Solve each of those separately for m.

OpenStudy (nekoaisu):

m = -5/4 or m = -3/2? @Directrix

Directrix (directrix):

Correct.

OpenStudy (nekoaisu):

@Directrix Thank You!

Directrix (directrix):

You are welcome.